# Language-centered Human Activity Recognition

### v3

@@ 3.6. Training and inference process
Inference process: For inference, the sensor encoder and activity label embeddings are deployed on mobile devices. IMU sensor readings are input into the encoder to generate embeddings, which are then compared with the activity label embeddings to calculate similarity. The activity label with the highest similarity is identified as the recognized activity. This process is also used for recognizing new activities.
@@ 4.1.2. Data preprocessing
We preprocess all datasets to ensure consistent sampling rates and window sizes. The sampling rate is set to 20 Hz to simulate a resource-constrained scenario, following prior studies (Xu et al., 2023). For cross-dataset activity recognition, we focus on four common activities (walking, going upstairs, going downstairs, and sitting) across the four datasets. We select one dataset as the source and one of the remaining three as the target, resulting in 12 combinations in total. Datasets are categorized based on their roles as either source or target. For source datasets, each is split into a training set (80%) and a validation set (20%). For target datasets, the entire dataset is used as the test set to directly assess cross-dataset performance. For new activity recognition, we train the model on the four common activities from the source dataset. We then select activities from the target dataset that differ from the four common activities as new activities. Finally, we evaluate the model’s ability to recognize these new activities by determining which activity labels the new data correspond to.
@@ 4.1.3. Baselines
(ii) Cross-dataset HAR baselines: SDMix (Lu et al., 2022) It is a method for cross-domain activity recognition that uses the mixup method that addresses semantic inconsistencies caused by domain gaps. UDAHAR (Chang et al., 2020) is a method for cross-dataset activity recognition using domain adaptation. UniHAR (Xu et al., 2023) is a method for cross-dataset activity recognition. It designs physical-informed augmentation methods to improve the cross-dataset HAR performance. UniMTS (Zhang et al., 2024) is a unified pre-training framework designed for motion time-series data.
@@ 4.1.5. Implementation
We implement our method and baselines using PyTorch 1.9.0 in a Python 3.7 environment and train them on two NVIDIA RTX A5000 GPUs, each with 24 GB of memory. For LLMs, we choose GPT-4 as our LLMs to obtain semantic interpretations of sensor readings and activity labels. For the text encoder, we use the BERT-base architecture as its structure and initialize it with the pre-trained BERT-base parameters (Devlin et al., 2018). For the sensor encoder, we set the embedding dimension to $768$ to maintain consistency with pre-trained BERT. The multi-head attention parameter is set to $2$ and the encoder layer consists of $3$ layers. We utilize the AdamW optimizer with a learning rate of 1e-5 and set the batch size to 32. For iterative re-generation, we set the threshold for intra-cluster distance change to 0.0005.
4.2. Overall performance
We compare our method on various baselines, and the results are presented in Table 2. Overall, our method demonstrates superior average performance across 12 different source and target dataset combinations, leading in the majority of settings. Compared to conventional HAR baselines, our method consistently outperforms them. These traditional baselines are not specifically designed for cross-dataset HAR, and their strong performance typically depends on large, carefully curated datasets. Compared to cross-dataset HAR baselines, our method outperforms them (except in the setting from Shoaib to Motion). Although SDMix and UDAHAR are also designed for cross-dataset HAR, they have specific requirements. For instance, UDAHAR may be more suitable for datasets with smaller distributional gaps. This consistent performance demonstrates the robustness and reliability of our method in diverse conditions and datasets.
@@ 4.2. Overall performance
UDAHAR Accuracy 0.473 0.481 0.249 0.453 0.406 0.448 0.453 0.406 0.448 0.323 0.285 0.533 0.413
F1 scores 0.317 0.407 0.191 0.306 0.254 0.302 0.306 0.254 0.302 0.248 0.165 0.489 0.298
@@ 4.6.1. The effect text encoder
Figure 11. Category-level HAR Performance
Figure 12. The Effect of Text Encoder
@@ 4.8. Deployment analysis
To evaluate the suitability of the sensor encoder for mobile deployment, we analyze its parameter size and inference time. Deploying LanHAR on mobile devices involves three key steps: (i) developing an Android application, (ii) converting the model into a .pt file, and (iii) integrating it for processing accelerometer and gyroscope data for activity recognition. For our case study, we select the Pixel 7 smartphone as the deployment platform. The sensor encoder, combined with the activity label embeddings, occupies approximately 40 MB of memory, of which 0.01 MB is allocated for the embeddings. To assess its real-time performance, we conduct 500 inference trials, measuring the time required to recognize an activity. The results indicate an average processing time of 0.027 seconds, with the minimum and maximum times recorded at 0.037 seconds and 0.023 seconds, respectively. For comparison, we evaluate the model’s inference efficiency on a single A5000 GPU with a batch size of 1. The average inference time for a single activity is 0.0007 seconds, with the minimum and maximum times recorded at 0.0006 seconds and 0.001 seconds, respectively. Although inference on the smartphone is significantly slower than on the GPU, it remains within a practical range for real-world mobile applications, supporting real-time inference with an IMU sampling frequency of around 40 Hz (i.e., 0.025 seconds per reading). This demonstrates that LanHAR can operate effectively on mobile devices. To further enhance efficiency, future optimizations, such as quantization and hardware acceleration, could be explored.
@@ 5. Discussion
Limitations: The process of obtaining semantic interpretations using LLMs is time-consuming, primarily due to the slow inference speed of current LLMs. We anticipate that future advancements in accelerating LLMs will enhance the efficiency of our method. Although we used four public datasets in the evaluation, they remain relatively small in scale with a limited number of activities. A larger dataset could provide a more comprehensive evaluation and further validate our method’s effectiveness.
6. Related work
Human activity recognition: Human Activity Recognition (HAR) leverages data from various sensors such as accelerometers and gyroscopes to detect and classify human activities (Yao et al., 2017). To mitigate the issue of data heterogeneity, existing works can be classified into three categories, including self-supervised learning methods, domain adaptation methods, and data augmentation methods. Self-supervised methods (Xu et al., 2021; Saeed et al., 2019) aim to maximize the use of unlabeled data to extract valuable features. This method allows models to require only a minimal amount of labeled data to efficiently adapt to downstream tasks. Despite their capabilities, these models still depend on a certain amount of labeled data to train classifiers, which may struggle when applied to target datasets that lack any labeled data. The core idea of domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) involves taking a classifier that has been pre-trained on a source domain and using it with an unlabeled dataset from a target domain, then adjusting the source model’s weights to enhance its performance in the target domain. Data augmentation (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019; Wang et al., 2022; Tang et al., 2021) in HAR aims to increase the diversity and quality of training data, thereby enhancing the model’s generalization ability and accuracy. However, these methods still face data heterogeneity and heterogeneity can amplify the impact of classification errors on downstream applications. Additionally, cross-dataset scenarios often introduce new activities that existing methods struggle to handle.
Some works (Cheng et al., 2013) attempts to address the challenge of recognizing new activities by using pre-defined semantics, such as actions of specific body parts. These methods train two mappings: one from sensor readings to predefined semantics, and another from predefined semantics to activity labels. However, the quality of these predefined semantics significantly influences the model performance. Additionally, this work is not well-suited for cross-dataset new activities, as the new activities may differ significantly from the original dataset.
LLM for human activity recognition: Recently, there has been a surge of interest in leveraging the capabilities of large language models (LLMs) to enhance mobile computing applications (Leng et al., 2024). UniHAR (Xu et al., 2023) explores the potential of leveraging large language models (LLMs) to comprehend and interact with the physical world. Kim et al. (Kim et al., 2024) harness the power of large language models (LLMs) to improve health prediction using data collected from wearable sensors. IMUGPT2.0 (Leng et al., 2024) utilizes LLMs to help data generation since IMU-labeled data is challenging to collect. MotionGPT (Jiang et al., 2023) fuses language data with large-scale motion models, and motion-language pre-training that can enhance the performance of motion-related tasks becomes feasible.
@@ 7. Conclusion
We design LanHAR, a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels for cross-dataset human activity recognition and new activity recognition. It adopts a two-stage training framework to transfer the capabilities of LLMs to mobile devices. The experimental results demonstrate that the proposed approach outperforms other state-of-the-art methods.

### v4

@@ 4.6. Training and inference process
Inference process: The model operates in an inference-only mode and does not require access to additional unlabeled data, as both the sensor encoder and activity label embeddings are pre-trained and deployed on the device. At inference time, IMU sensor readings are encoded and compared against the pre-stored activity label embeddings to compute similarity scores. The label with the highest similarity is selected as the recognized activity. This same process also applies when recognizing new activities.
@@ 5.1.2. Data preprocessing
We preprocess all datasets to ensure consistent sampling rates and window sizes. The sampling rate is set to 50 Hz to simulate a resource-constrained scenario, following prior studies (Xu et al., 2023). For cross-dataset activity recognition, we focus on four common activities (walking, going upstairs, going downstairs, and still) that are shared across the four datasets. Following prior work (Xu et al., 2023), we group sitting and standing under the unified category still. We select one dataset as the labeled source and one of the remaining three as the target, resulting in 12 source–target combinations. Datasets are categorized based on their roles as either source or target. For source datasets, each is split into a training set (80%) and a validation set (20%). For target datasets, the entire dataset is used as the test set to directly assess cross-dataset performance. In addition, during the second stage of training, we incorporate unlabeled data from the remaining source datasets to jointly train the sensor encoder. This helps the model learn more generalizable representations, even though the main supervision still comes from a single labeled source. For new activity recognition, we train the model on the four common activities from the source dataset. We then select activities from the target dataset that differ from the four common activities as new activities. Finally, we evaluate the model’s ability to recognize these new activities by determining which activity labels the new data correspond to.
@@ 5.1.3. Baselines
(ii) Cross-dataset HAR baselines: SDMix (Lu et al., 2022) It is a method for cross-domain activity recognition that uses the mixup method that addresses semantic inconsistencies caused by domain gaps. UDAHAR (Chang et al., 2020) In our setting, we follow the feature alignment methods UDA paradigm, which aims to minimize the domain gap by explicitly aligning the feature distributions of the source and target domains. UniHAR (Xu et al., 2023) is a method for cross-dataset activity recognition. It designs physical-informed augmentation methods to improve the cross-dataset HAR performance. UniMTS (Zhang et al., 2024) is a unified pre-training framework designed for motion time-series data.
@@ 5.1.5. Implementation
We implement our method and baselines using PyTorch 1.9.0 in a Python 3.7 environment and train them on two NVIDIA RTX A5000 GPUs, each with 24 GB of memory. For LLMs, we choose GPT-4 as our LLMs to obtain semantic interpretations of sensor readings and activity labels. For the text encoder, we use the BERT architecture as its structure and initialize it with the pre-trained BERT parameters (Devlin et al., 2018). For the sensor encoder, we set the embedding dimension to $768$ to maintain consistency with pre-trained BERT. The multi-head attention parameter is set to $2$ and the encoder layer consists of $3$ layers. We utilize the AdamW optimizer with a learning rate of 1e-5 and set the batch size to 32. For iterative re-generation, we set the threshold for intra-cluster distance change to 0.0005.
@@
UDAHAR Accuracy 0.473 0.481 0.249 0.446 0.417 0.315 0.428 0.559 0.591 0.452 0.551 0.561 0.460
F1 scores 0.317 0.407 0.191 0.279 0.336 0.292 0.355 0.471 0.473 0.265 0.398 0.281 0.339
@@
5.2. Overall performance
5.2.1. Comparison to baselines using single labeled source dataset.
We compare our method on various baselines, and the results are presented in Table 2. Overall, our method demonstrates superior average performance across 12 different source and target dataset combinations, leading in the majority of settings. Compared to conventional HAR baselines, our method consistently outperforms them. These traditional baselines are not specifically designed for cross-dataset HAR, and their strong performance typically depends on large, carefully curated datasets. Compared to cross-dataset HAR baselines, our method outperforms them (except in the setting from Shoaib to Motion). Although SDMix and UDAHAR are also designed for cross-dataset HAR, they have specific requirements. For instance, UDAHAR may be more suitable for datasets with smaller distributional gaps. This consistent performance demonstrates the robustness and reliability of our method in diverse datasets.
Following the reviewer’s suggestion, we extend our cross-dataset experiments to include all relevant activities, treating sitting and standing as separate classes rather than merging them into a single still category. Specifically, the evaluated activities now include sitting, standing, walking, going upstairs, and going downstairs. The new cross-dataset results for five activities are as shown in Appendix Table 5. As shown in the results, the inclusion of both sitting and standing as separate activities introduces additional difficulty, as these two activities exhibit highly similar sensor patterns and are often hard to distinguish, especially in cross-dataset settings. This leads to a noticeable performance drop across all methods. While the performance of our method also drops due to the increased difficulty, it still achieves the best results among all compared approaches. This suggests that, despite the challenge of distinguishing highly similar activities such as sitting and standing, our method remains comparatively more effective.
5.2.2. Comparison to baselines using multiple labeled source datasets.
In addition to baselines trained on a single labeled source dataset, some existing works aim to mitigate cross-dataset heterogeneity by leveraging multiple labeled source datasets. In this section, we compare our method with two types of approaches. The specific implementations are as follows: (i) GOAT (Miao and Chen, 2024): We implemented the GOAT framework following the original paper’s setting, where the model is pre-trained on three datasets and evaluated on a separate target dataset. Since most public datasets do not provide sensor position annotations, we removed the device location descriptions to ensure consistent input formats across datasets. (ii) Combining Public HAR Datasets (Presotto et al., 2023): We implement the two base models used in this work (i.e., HART (Ek et al., 2022) and DeepConvLSTM (Ordóñez and Roggen, 2016)) and adopt the multi-source pre-training setup on three datasets. However, the original method assumes access to the target dataset for fine-tuning, while our method does not require any target-domain labels. To ensure a fair comparison, we evaluated their approach under two settings: (i) No fine-tuning: testing on the target dataset without fine-tuning, and (ii) Fine-tuning: using 5% of labeled target data. Results are reported in Table 3.
From the results in the no fine-tuning scenario, we observe that merely using more data for pre-training does not guarantee improved performance. While aggregating multiple datasets can help reduce domain differences, effective performance typically requires an extensive and diverse set of source data. Otherwise, substantial distribution gaps between source and target domains still limit the generalization of pre-trained models. In contrast, although fine-tuning significantly boosts performance, it relies heavily on labeled target-domain data, which is frequently unavailable in practical scenarios. Our proposed method achieves robust performance without requiring any labeled data from the target domain. This demonstrates superior adaptability, performing notably better than other methods without fine-tuning and comparably to methods that use fine-tuning.
Table 3. Comparison to baselines using multiple labeled source datasets. Bolded scores indicate the best.
Method Metric HHAR Motion UCI Shoaib
Goat Accuracy 0.431 0.549 0.704 0.535
F1 scores 0.156 0.262 0.573 0.402
Public-HART (No fine-tuning) Accuracy 0.432 0.542 0.695 0.593
F1 scores 0.169 0.253 0.567 0.462
Public-DeepConv (No fine-tuning) Accuracy 0.458 0.576 0.731 0.646
F1 scores 0.221 0.311 0.609 0.545
Public-HART (Fine-tuning) Accuracy 0.798 0.803 0.830 0.862
F1 scores 0.731 0.764 0.779 0.830
Public-DeepConv (Fine-tuning) Accuracy 0.792 0.762 0.822 0.767
F1 scores 0.673 0.697 0.757 0.703
LanHAR Accuracy 0.756 0.797 0.794 0.747
F1 scores 0.704 0.760 0.749 0.712
@@ 5.4. Complex activity performance
5.5. Non-cross-dataset performance
To evaluate the model’s performance in a standard within-dataset setting, we conduct additional non-cross-dataset experiments. In this setup, the model is trained and tested on the same dataset (i.e., HHAR, Motion, UCI, and Shoaib). Specifically, each dataset is split into training and testing sets using an 8:2 ratio. The results are as shown in Table 4.
Table 4. Non-cross-dataset performance. Bolded scores indicate the best.
Method Metric HHAR UCI Motion Shoaib
DCNN Accuracy 0.995 0.947 0.980 0.982
F1 scores 0.993 0.954 0.967 0.982
Transformer Accuracy 0.932 0.924 0.926 0.940
F1 scores 0.929 0.918 0.903 0.935
LIMU-BERT Accuracy 0.961 0.976 0.983 0.940
F1 scores 0.961 0.962 0.980 0.947
UDAHAR Accuracy 0.992 0.930 0.980 0.978
F1 scores 0.991 0.938 0.967 0.978
SDMix Accuracy 0.949 0.957 0.983 0.924
F1 scores 0.949 0.934 0.982 0.931
UniMTS Accuracy 0.936 0.935 0.953 0.961
F1 scores 0.912 0.924 0.947 0.954
UniHAR Accuracy 0.927 0.937 0.971 0.900
F1 scores 0.886 0.904 0.949 0.882
LanHAR Accuracy 0.956 0.951 0.922 0.967
F1 scores 0.949 0.954 0.906 0.965
As shown in the results, most methods, including ours, achieve high accuracy in this setting, likely because each single dataset is collected from a relatively homogeneous group of participants under similar conditions, resulting in low distributional variability. Although our model is not the best-performing method on every dataset, it remains highly competitive. In terms of accuracy, the best-performing method on each dataset outperforms the worst-performing one by an average of 7.17%, while the gap between the best-performing method and ours is only 3.72% on average. Similarly, for F1-score, the average difference between the best and worst methods per dataset is 9.05%, while the best-performing method exceeds ours by just 3.91% on average. The slightly lower performance of our model may be due to the loss of some dataset-specific information during the semantic interpretation process, which helps generalization across datasets but may limit performance within a single dataset.
@@ 5.6. Category-level HAR performance
Figure 11. Category-level HAR Performance
Figure 12. The Effect of Text Encoder
@@ 5.9. Deployment analysis
To evaluate the suitability of the sensor encoder for mobile deployment, we analyze its parameter size and inference time. Deploying LanHAR on mobile devices involves three key steps: (i) developing an Android application, (ii) converting the model into a .pt file, and (iii) integrating it for processing accelerometer and gyroscope data for activity recognition. For our case study, we select the Pixel 7 smartphone as the deployment platform. The sensor encoder, combined with the activity label embeddings, occupies approximately 40 MB of memory, of which 0.01 MB is allocated for the embeddings. To assess its real-time performance, we conduct 100 inference trials, measuring the time required to recognize an activity. The results indicate an average processing time of 0.607 seconds, with the minimum and maximum times recorded at 0.716 seconds and 0.527 seconds, respectively. For comparison, we evaluate the model’s inference efficiency on a single A5000 GPU with a batch size of 1. The average inference time for a single activity is 0.0007 seconds, with the minimum and maximum times recorded at 0.0006 seconds and 0.001 seconds, respectively. Although inference on the smartphone is significantly slower than on the GPU, it remains within a practical range for real-world mobile applications, supporting real-time inference with an IMU sampling frequency of around 40 Hz (i.e., 0.025 seconds per reading). This demonstrates that LanHAR can operate effectively on mobile devices. To further enhance efficiency, future optimizations, such as quantization and hardware acceleration, could be explored.
@@ 6. Discussion
Limitations: (i) The process of obtaining semantic interpretations using LLMs is time-consuming, primarily due to the slow inference speed of current LLMs. We anticipate that future advancements in accelerating LLMs will enhance the efficiency of our method. (ii) Although we used four public datasets in the evaluation, they remain relatively small in scale with a limited number of activities. A larger dataset could provide a more comprehensive evaluation and further validate our method’s effectiveness. (iii) Although our evaluation on PAMAP2 simulates domain heterogeneity by splitting participants, it does not constitute a true cross-dataset validation. This design choice was motivated by the need to assess model performance on complex activities (e.g., vacuum cleaning, ironing), which are sparsely available and inconsistently defined across public datasets. Even though, we believe this setting is usable to evaluate the performance on complex activities when no additional datasets with the same complex activities are available.
@@ 7. Conclusion
In this work, we incorporate LLM-generated semantic interpretations of both sensor readings and activity labels to mitigate cross-dataset heterogeneity and to enable the recognition of unseen activities by mapping both sensor readings and activity labels into a shared semantic space. We design LanHAR, a lightweight system that features a semantic interpretation generation process with an iterative re-generation mechanism to ensure output quality. Our two-stage training and inference framework allows the transfer of LLM capabilities to mobile and resource-constrained environments, enabling practical deployment scenarios. We conduct comprehensive experiments across five public HAR datasets. The results demonstrate that our method consistently outperforms state-of-the-art approaches in cross-dataset activity recognition. In future work, we plan to extend our framework to support open-set activity recognition, integrate physics-informed semantic interpretations and handle more complex, fine-grained activity patterns to further improve robustness and generalizability.
Appendix
Comparison to baselines with 5 simple activities
Following the reviewer’s suggestion, we extend our cross-dataset experiments to include all relevant activities, treating sitting and standing as separate classes rather than merging them into a single still category. Specifically, the evaluated activities now include sitting, standing, walking, going upstairs, and going downstairs. The new cross-dataset results for five activities are as shown in Table 5.
Table 5. Overall performance with 5 simple activities (walking, going upstairs, going downstairs,sitting and standing). Bolded scores indicate the best, while underlined scores represent the second best.
Method Source HHAR UCI Motion Shoaib Average
DCNN Accuracy 0.231 0.154 0.181 0.271 0.175 0.292 0.245 0.342 0.518 0.323 0.462 0.362 0.296
F1 scores 0.162 0.091 0.084 0.163 0.107 0.251 0.092 0.246 0.486 0.208 0.405 0.218 0.209
Transformer Accuracy 0.285 0.266 0.270 0.230 0.244 0.213 0.221 0.246 0.441 0.220 0.372 0.277 0.274
F1 scores 0.195 0.198 0.205 0.078 0.158 0.216 0.072 0.182 0.416 0.072 0.287 0.105 0.182
LIMU-BERT Accuracy 0.264 0.214 0.408 0.411 0.230 0.262 0.186 0.129 0.275 0.442 0.313 0.224 0.279
F1 scores 0.150 0.139 0.350 0.381 0.075 0.191 0.076 0.064 0.165 0.412 0.234 0.073 0.193
UDAHAR Accuracy 0.266 0.316 0.348 0.270 0.173 0.376 0.228 0.286 0.468 0.254 0.496 0.304 0.315
F1 scores 0.128 0.215 0.250 0.125 0.104 0.339 0.074 0.206 0.447 0.137 0.436 0.155 0.218
SDMix Accuracy 0.375 0.349 0.433 0.462 0.390 0.461 0.371 0.327 0.300 0.418 0.197 0.518 0.383
F1 scores 0.244 0.273 0.370 0.395 0.245 0.441 0.300 0.281 0.172 0.386 0.135 0.435 0.306
UniMTS Accuracy 0.439 0.457 0.482 0.441 0.402 0.511 0.407 0.437 0.514 0.492 0.402 0.553 0.461
F1 scores 0.355 0.361 0.383 0.392 0.328 0.405 0.301 0.382 0.416 0.354 0.327 0.468 0.372
UniHAR Accuracy 0.575 0.523 0.489 0.475 0.569 0.628 0.553 0.570 0.551 0.464 0.569 0.579 0.545
F1 scores 0.470 0.399 0.413 0.421 0.489 0.586 0.457 0.510 0.480 0.397 0.491 0.501 0.467
LanHAR Accuracy 0.629 0.585 0.557 0.576 0.624 0.631 0.587 0.611 0.558 0.561 0.596 0.580 0.591
F1 scores 0.510 0.461 0.447 0.443 0.528 0.588 0.502 0.527 0.484 0.492 0.512 0.505 0.497